Circular No. 8770

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006fo AND 2006le–2006lu Further to IAUC 8766, seven apparent supernovae have been reported

from unfiltered CCD images: 2006le, 2006lf, 2006lg, 2006li, and 2006lu by R. R. Prasad, N. Joubert, M. Baek, and W. Li (LOSS/KAIT); and 2006lh and 2006lt by the “Nearby Supernova Factory” collaboration (details on CBETs 708 and 714).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006le Oct. 26.44 5^{h}00^{m}41.^{s}99 +62^{o}15′19′′.0 18.0 12′′.4 W, 40′′.1 N
2006lf Oct. 26.43 4 38 29.49 +44 02 01.5 18.0 13′′.0 E, 12′′.1 S
2006lg Oct. 25.38 2 46 57.64 +39 00 40.0 18.8 0′′.6 W, 4′′.8 S
2006lh Oct. 24.2 23 26 51.60 + 9 19 57.6 19.4  –
2006li Oct. 28.49 7 05 33.33 +33 50 01.3 19.4 3′′.3 W, 20′′.4 N
2006lt Oct. 11.4 2 16 59.89 +30 41 57.4 19.1 –
2006lu Oct. 30.55 9 15 17.63 −25 36 00.3 17.6 6′′.0 E, 0′′.5 S

Additional KAIT magnitudes for 2006le in UGC 3218: Oct. 19.44 UT, [19.1; 25.45, 18.5; 27.51, 17.6. SN 2006le is a type-Ia supernova, well before maximum light on Oct. 27 (details on CBET 702). Additional KAIT magnitudes for 2006lf in UGC 3108: Oct. 19.43, [19.3; 27.51, 17.6. SN 2006lf is a type-Ia supernova, around 1 week before maximum light on Oct. 28 (details on CBET 705). Additional KAIT magnitudes for 2006lg in UGC 2243: Oct. 1.38, [19.0; 26.36, 18.4; 28.34, 18.5. SN 2006lg is a type-II supernova (details on CBET 707). A spectrogram obtained on Oct. 28 showed 2006lh to be a young type-II supernova (details on CBET 708). Additional magnitudes for 2006li in UGC 3664: 1989 Oct. 28, [21.0 (Digitized Sky Survey red plate); 2003 Oct. 22.52, 19.8 (KAIT); 2006 Oct. 29.52, 19.4 (KAIT). A spectrogram obtained on Oct. 13 suggests that 2006lt was then already a late type-Ib supernova. Additional magnitudes for 2006lu: May 4.12, [18.8 (KAIT); Nov. 1.35, r = 17.51 ± 0.03 (W. Krzeminsky and C. Contreras, on behalf of the Carnegie Supernova Project; Las Campanas Observatory 1-m Swope telescope).

J. Frieman, on behalf of the Sloan Digital Sky Survey II collaboration, has further reported ten additional faint supernovae (2006lj–2006ls; magnitude range: g = 21.3–23.5); details are given on CBET 713. Frieman’s group has also noted that further analysis of the spectrum of 2006fo (cf. IAUC 8750) shows it to be consistent with a type-Ic supernova, with features similar to that of 1999ex (details on CBET 643).

2006 November 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green